OpenStudy (anonymous):

Write each fraction or mixed number as decimal. 1 1/4

OpenStudy (anonymous):

1 is a whole number

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

1 1/4 = 5/4 125/100=1.25